OpenStudy (anonymous):

Help with algebra 2 question please? Evaluate log2 (1/8). I know the answer is -3....I just don't know how to get the answer.

OpenStudy (anonymous):

log2 (1/8) is basically asking: 2^(x)=1/8 Well 2^3=8 So if you make the exponent negative, you'll get 1/8 So that's why the answer is -3 :) Hope that helps!

OpenStudy (mathstudent55):

\(\log_b b^n = n\) Write 1/8 as a power of 2. Then the log of 1/8 is simply the exponent that 2 is raised to.

OpenStudy (anonymous):

Bruhhh thank you so much. :)

OpenStudy (anonymous):

No prob! :)